Area Between Curves, Volumes and Average Values of Functions:


1 Calculate the area between the parabola 𝑦=π‘₯2 and the line 𝑦=4


2 Derive the formula for the volume of the right-circular cone using a single integral given that 𝑦=βˆ’π‘Žβ„Ž(π‘§βˆ’β„Ž).



3. Find the average value 𝑓(π‘₯)Μ…Μ…Μ…Μ…Μ…Μ… of 𝑓(π‘₯)=2π‘₯2+3π‘₯+3 in the interval [1,4].



1
Expert's answer
2021-11-22T18:21:27-0500

1.


"x^2=4=>x_1=-2, x_2=2"

"Area=A=\displaystyle\int_{-2}^{2}(4-x^2)dx=[4x-\dfrac{x^3}{3}]\begin{matrix}
 2 \\
 -2
\end{matrix}"

"=8-\dfrac{8}{3}-(-8+\dfrac{8\dfrac{a}{b}}{3})=\dfrac{32}{3}({units}^2)"

2.

"𝑦=−\dfrac{a}{h}(𝑧−ℎ)"

"V=\displaystyle\int_{0}^{h}\pi(y(z))^2dz=\displaystyle\int_{0}^{h}\pi(−\dfrac{a}{h}(z-h))^2dz"

"=\pi \dfrac{a^2}{h^2}[\dfrac{(z-h)^3}{3}]\begin{matrix}
 h \\
 0
\end{matrix}=\pi \dfrac{a^2}{3h^2}(0+h^3)=\dfrac{1}{3}\pi a^2 h({units}^3)"

3.


"f_{ave}=\dfrac{1}{4-1}=\displaystyle\int_{1}^{4}(2x^2+3x+3)dx"

"=\dfrac{1}{3}[\dfrac{2x^3}{3}+\dfrac{3x^2}{2}+3x]\begin{matrix}
 4 \\
 1
\end{matrix}"

"=\dfrac{1}{3}(\dfrac{128}{3}+24+12-\dfrac{2}{3}-\dfrac{3}{2}-3)"

"=\dfrac{49}{2}"